cg188290:

The maximum height of a vehicle that can safely pass under a bridge is 12 feet 5 inches. A truck measures 162 inches in height. Which best explains whether or not the truck can pass safely under the bridge?

Vocaloid:

Answer choices A. The truck can pass safely under the bridge. The truck is 13 inches shorter than the maximum height. B. The truck cannot pass safely under the bridge. The truck is 13 inches taller than the maximum height. C. The truck can pass safely under the bridge. The truck is 12 inches shorter than the maximum height. D. The truck cannot pass safely under the bridge. The truck is 12 inches taller than the maximum height

Answermyquestions:

transfer the inches into feet.Remember that 12 inches = 1 foot

mhanifa:

12 feet and 5 inches = 12*12 + 5 inches = 149 inches Compare the height of the bridge with the height of the truck: 162 - 149 = 13 inches Correct choice is B.